# An Incomplete Loop: Deductive, Inductive, and Abductive Learning in Large Language Models

### v1

@@ 2 Three Types of Reasoning in Language Models
Reasoning type Analogue in LMs Citations
Deductive Instruction following Wei et al. (2021); Sanh et al. (2021)
Inductive In-context learning Brown et al. (2020)
Abductive Chain-of-thought (one hypothesis) Wei et al. (2023); Kojima et al. (2022)
Abductive Externally aided instruction inference (many hypotheses) Qiu et al. (2023); Wang et al. (2023); Zhu et al. (2023); Yang et al. (2024),
Abductive LM-guided instruction inference (many hypotheses) This paper
@@ 4 Domains and Evaluation
Domain In-context examples ($D_{i}$) Query example ($x_{i}$) Expected answer ($y_{i}$) Hypothesis example ($h_{i}$)
Functions (-10, -213), (9, 167), (4, 67), … 15 287 $f(x)=20x-13$
Colours (lug dax, blue green), (lug zup, blue yellow), (lug bluf, blue blue), … lug walm dax bluf blue blue blue green green lug $\rightarrow$ blue
Kalamang vocab (ek) (That is their place., Tompat ma me muin), (I’m getting pandandus, I want to make a mat.), … Sakina is pouching guavas. Sakina sarimara lawat. guava $\rightarrow$ sarim
Kalamang grammar (ek) “” Order of subject and verb: SV
@@ 5 Results
Domain Deductive reasoning works on average Abductive reasoning works on average Hypothesis proposal works on average Abductive reasoning related to inductive reasoning
@@ Appendix B Prompts for Linear Functions Domain
Prompt Type Usage Prompt Text
@@
Base system prompt For reasoning with in-context examples You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.
@@
Hypothesis proposal system prompt For proposing hypotheses based on in-context examples You are a pattern recognition system. Your job is to come up with a function that describes the data as well as you can.
@@
Instruction following system prompt For applying a proposed hypothesis or ground-truth hypothesis to the input You are a problem solving system. Your job is to apply the function to the data in order to produce an answer.
Few-shot examples prompt For reasoning with in-context examples only Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input}
@@
Prompt with ground-truth hypothesis Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included. Function: {The real function} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for hypothesis induction Used to have the model propose a single hypothesis for the function based on in-context examples Write the function that captures the relationship between inputs and outputs. You should write it in the form y = ax^0 + bx^1. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function (please write explicitly in the exact form" ‘Output: y = ax^0 + bx^1)’:
@@
Prompt with a self-induced hypothesis Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed. Use this function to apply to the input example to get the correct output. { model’s hypothesis } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for zero-shot chain-of-thought Used to encourage the model to generate a chain of thought. Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the function could be. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
@@
Prompt for hypothesis probability estimate Used to prompt a language model directly for reranking hypotheses How likely is this hypothesis about the function to be true given the data? Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function explanation: {model’s hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
@@
Prompt for data logprobs given hyp estimate Used to rerank hypotheses based on logprobs. These are examples of applying this function: {model’s hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...
@@ Appendix F Prompts for Colours Domain
Prompt Type Usage Prompt Text
@@
Base system prompt For reasoning with in-context examples You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.’’
@@
Hypothesis proposal system prompt For proposing hypotheses based on in-context examples You are a rule induction system. Your job is to figure out the rules underlying a problem and report on them. Use the examples to guide your thinking.
@@
Instruction following system prompt For applying a proposed hypothesis or ground-truth hypothesis to the input You are a parser. Carefully use the grammar to parse inputs to determine the correct output.
@@
Few-shot examples prompt For reasoning with in-context examples only Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input} Remember to start your answer with ’Output:’
@@
Prompt with ground-truth hypothesis Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included. Use this grammar to parse the input example to get the correct output. Grammar: {colours domain grammar} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for hypothesis induction Used to have the model propose a single hypothesis for the translation of a word. The below examples contain the nonce word {word}. Using the examples, deduce what {word} means. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Write your answer like this: word -> meaning. Meaning can be a word or a general rule dependent on the context. Rule:
@@
Prompt with a self-induced hypothesis Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed. Use this grammar to parse the input example to get the correct output. { model’s hypothesis grammar } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for zero-shot chain-of-thought Used to encourage the model to generate a chain of thought. Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the translation could be. Work through your answer step by step and show your work. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
@@
Prompt for hypothesis probability estimate Used to prompt a language model directly for reranking hypotheses These are examples of the translation of the word {word}. Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
@@
Prompt for data logprobs given hyp estimate Used to rerank hypotheses based on logprobs. These are examples of applying this function: {model’s word hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...
@@ Appendix I MTOB Experimental Settings
Prompt Type Usage Prompt Text
Base system prompt For reasoning with in-context examples ‘‘’’
Hypothesis proposal system prompt (Grammar) For proposing grammar feature hypotheses based on in-context examples ‘‘’’
Hypothesis proposal system prompt (Vocab) For proposing vocabulary mappings based on in-context examples ‘‘’’
Instruction following system prompt For applying a proposed hypothesis or ground-truth hypothesis to the input ‘‘’’
@@
Few-shot examples prompt For reasoning with in-context examples only Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Now write the translation. English: {query input} Kalamang translation:
Prompt with ground-truth hypothesis Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included. Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {grammar sketch} Now write the translation. English: {query input} Kalamang translation:
Prompt for hypothesis induction Used to have the model propose a single hypothesis for the translation of a word. The following sentences contain the word ‘{word}’ in English. Examples: English sentence: {eng sentence} Kalamang translation: {kgv sentence} ... What is the Kalamang translation of the word ‘{word}’? Write your answer like this: {English word} -> {Kalamang translation}.
@@
Prompt with a self-induced hypothesis Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed. Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{hypothesis translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {hypothesis grammar sketch} Now write the translation. English: {query input} Kalamang translation:
@@
Prompt for zero-shot chain-of-thought Used to encourage the model to generate a chain of thought. Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Let’s think step by step about what the translation could be. Now write the translation. English: {query input} Kalamang translation:
@@
Prompt for hypothesis probability estimate Used to prompt a language model directly for reranking hypotheses These are examples of the translation of the word {word}. Examples: {example translation pairs} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
@@
Prompt for data logprobs given hyp estimate Used to rerank hypotheses based on logprobs. These are examples of sentences that contain the word {word}: This is the translation of the word: {model’s word hypothesis} Examples: {example translation pairs}
@@
Source Reference Translation (few-shot) Translation (instruction-induction:verbalized confidence) Translation (true-instruction)
Who asked you? Namana kat gerket? Naman mua kiemara kowam? Naman era? Naman mua
@@
The rope is wrapped around it. Karek mat kokarap. Kainasu esun kiwat semda bo dalangdalang. Sar tok esun et dalangdalang na. Karek mambon guanggarien ma he me.
Before, they said Japan and The Netherlands were at war. Wiseme mu toni Japangbon Beladarbon nausair. Wiseme mu Jafaris emun Nabil emun kaden kiesun ning. Wise me mu salai Japan an Sar Beladargara din jambat mara. Tikninda, termus siram panci tubak Beladar tamatko pang.
Mu bo Tamisengga som dande yecie. They are going (on a return trip) to bury someone in Antalisa. They go to Tamisengga and return. They go to Antalisa, someone died, and they return. They will return to Tamisengga until they bury the person.
Suo me kene arkon metko. At Suo is a kene tree. I washed my body in the river. Back me gives birth canoe. Wash your face near the Narkon tree.
An istrat kahendengoa marmar ba temun. I’m walking on a long wide road. I put the long stick on the ground. I have a long drying rack but I am walking. I walked far down the street to my uncle.

### v2

@@ 2 Three Types of Reasoning in Language Models
Reasoning type Analogue in LMs
Citations
Deductive Instruction following
Wei et al. (2021); Sanh et al. (2021)
Inductive In-context learning
Brown et al. (2020)
Abductive Chain-of-thought (one hypothesis)
Wei et al. (2023); Kojima et al. (2022)
Abductive Externally aided instruction inference (many hypotheses)
Qiu et al. (2023); Wang et al. (2023); Zhu et al. (2023); Yang et al. (2024),
Abductive LM-guided instruction inference (many hypotheses)
This paper
@@ 4 Domains and Evaluation
Domain
In-context examples ($D_{i}$)
Query example ($x_{i}$)
Expected answer ($y_{i}$)
Hypothesis example ($h_{i}$)
Functions
(-10, -213), (9, 167), (4, 67), …
15
287
$f(x)=20x-13$
Colours
(lug dax, blue green), (lug zup, blue yellow), (lug bluf, blue blue), …
lug walm dax bluf
blue blue blue green green
lug $\rightarrow$ blue
Kalamang vocab (ek)
(That is their place., Tompat ma me muin), (I’m getting pandandus, I want to make a mat.), …
Sakina is pouching guavas.
Sakina sarimara lawat.
guava $\rightarrow$ sarim
Kalamang grammar (ek)
Order of subject and verb: SV
@@ 5 Results
Domain
Deductive reasoning works on average
Abductive reasoning works on average
Hypothesis proposal works on average
Abductive reasoning related to inductive reasoning
@@ Appendix B Prompts for Linear Functions Domain
Prompt Type
Usage
Prompt Text
Base system prompt
For reasoning with in-context examples
@@
You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.
Hypothesis proposal system prompt
For proposing hypotheses based on in-context examples
@@
You are a pattern recognition system. Your job is to come up with a function that describes the data as well as you can.
Instruction following system prompt
For applying a proposed hypothesis or ground-truth hypothesis to the input
@@
You are a problem solving system. Your job is to apply the function to the data in order to produce an answer.
Few-shot examples prompt
For reasoning with in-context examples only
Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input}
Prompt with ground-truth hypothesis
@@
Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included.
Function: {The real function} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for hypothesis induction
Used to have the model propose a single hypothesis for the function based on in-context examples
@@
Write the function that captures the relationship between inputs and outputs. You should write it in the form y = ax^0 + bx^1. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function (please write explicitly in the exact form" ‘Output: y = ax^0 + bx^1)’:
Prompt with a self-induced hypothesis
Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed.
@@
Use this function to apply to the input example to get the correct output. { model’s hypothesis } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for zero-shot chain-of-thought
Used to encourage the model to generate a chain of thought.
@@
Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the function could be. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
Prompt for hypothesis probability estimate
Used to prompt a language model directly for reranking hypotheses
@@
How likely is this hypothesis about the function to be true given the data? Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function explanation: {model’s hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
Prompt for data logprobs given hyp estimate
Used to rerank hypotheses based on logprobs.
@@
These are examples of applying this function: {model’s hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...
@@ Appendix F Prompts for Colours Domain
Prompt Type
Usage
Prompt Text
Base system prompt
For reasoning with in-context examples
@@
You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.’’
Hypothesis proposal system prompt
For proposing hypotheses based on in-context examples
@@
You are a rule induction system. Your job is to figure out the rules underlying a problem and report on them. Use the examples to guide your thinking.
Instruction following system prompt
For applying a proposed hypothesis or ground-truth hypothesis to the input
@@
You are a parser. Carefully use the grammar to parse inputs to determine the correct output.
Few-shot examples prompt
For reasoning with in-context examples only
@@
Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input} Remember to start your answer with ’Output:’
Prompt with ground-truth hypothesis
@@
Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included.
Use this grammar to parse the input example to get the correct output. Grammar: {colours domain grammar} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for hypothesis induction
Used to have the model propose a single hypothesis for the translation of a word.
@@
The below examples contain the nonce word {word}. Using the examples, deduce what {word} means. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Write your answer like this: word -> meaning. Meaning can be a word or a general rule dependent on the context. Rule:
Prompt with a self-induced hypothesis
Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed.
@@
Use this grammar to parse the input example to get the correct output. { model’s hypothesis grammar } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for zero-shot chain-of-thought
Used to encourage the model to generate a chain of thought.
@@
Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the translation could be. Work through your answer step by step and show your work. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
Prompt for hypothesis probability estimate
Used to prompt a language model directly for reranking hypotheses
@@
These are examples of the translation of the word {word}. Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
Prompt for data logprobs given hyp estimate
Used to rerank hypotheses based on logprobs.
@@
These are examples of applying this function: {model’s word hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...
@@ Appendix I MTOB Experimental Settings
Prompt Type
Usage
Prompt Text
Base system prompt
For reasoning with in-context examples
Hypothesis proposal system prompt (Grammar)
For proposing grammar feature hypotheses based on in-context examples
Hypothesis proposal system prompt (Vocab)
For proposing vocabulary mappings based on in-context examples
Instruction following system prompt
For applying a proposed hypothesis or ground-truth hypothesis to the input
Few-shot examples prompt
For reasoning with in-context examples only
@@
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Now write the translation. English: {query input} Kalamang translation:
Prompt with ground-truth hypothesis
Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included.
@@
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {grammar sketch} Now write the translation. English: {query input} Kalamang translation:
Prompt for hypothesis induction
Used to have the model propose a single hypothesis for the translation of a word.
The following sentences contain the word ‘{word}’ in English. Examples: English sentence: {eng sentence} Kalamang translation: {kgv sentence} ... What is the Kalamang translation of the word ‘{word}’? Write your answer like this: {English word} -> {Kalamang translation}.
Prompt with a self-induced hypothesis
Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed.
@@
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{hypothesis translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {hypothesis grammar sketch} Now write the translation. English: {query input} Kalamang translation:
Prompt for zero-shot chain-of-thought
Used to encourage the model to generate a chain of thought.
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Let’s think step by step about what the translation could be. Now write the translation. English: {query input} Kalamang translation:
Prompt for hypothesis probability estimate
Used to prompt a language model directly for reranking hypotheses
@@
These are examples of the translation of the word {word}. Examples: {example translation pairs} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
Prompt for data logprobs given hyp estimate
Used to rerank hypotheses based on logprobs.
@@
These are examples of sentences that contain the word {word}: This is the translation of the word: {model’s word hypothesis} Examples: {example translation pairs}
@@
Source
Reference
Translation (few-shot)
Translation (instruction-induction:verbalized confidence)
Translation (true-instruction)
Who asked you?
Namana kat gerket?
Naman mua kiemara kowam?
Naman era?
Naman mua
@@
The rope is wrapped around it.
Karek mat kokarap.
Kainasu esun kiwat semda bo dalangdalang.
Sar tok esun et dalangdalang na.
Karek mambon guanggarien ma he me.
Before, they said Japan and The Netherlands were at war.
Wiseme mu toni Japangbon Beladarbon nausair.
Wiseme mu Jafaris emun Nabil emun kaden kiesun ning.
Wise me mu salai Japan an Sar Beladargara din jambat mara.
Tikninda, termus siram panci tubak Beladar tamatko pang.
Mu bo Tamisengga som dande yecie.
They are going (on a return trip) to bury someone in Antalisa.
They go to Tamisengga and return.
They go to Antalisa, someone died, and they return.
They will return to Tamisengga until they bury the person.
Suo me kene arkon metko.
At Suo is a kene tree.
I washed my body in the river.
Back me gives birth canoe.
Wash your face near the Narkon tree.
An istrat kahendengoa marmar ba temun.
I’m walking on a long wide road.
I put the long stick on the ground.
I have a long drying rack but I am walking.
I walked far down the street to my uncle.